OpenStudy (anonymous):

solve |5x – 2| 8 this ones a little trickier..

OpenStudy (lgbasallote):

there's really no equal signs at all?

OpenStudy (anonymous):

nope just greater than or equal tooo this might be a challenge for u @melbel

OpenStudy (anonymous):

is it like |5x-2| = 8 ??

OpenStudy (anonymous):

|5x-2| (greater than or equal to 8 sorry forgot to put that before

OpenStudy (anonymous):

take two conditions |5x-2| >= 8 |5x-2| >= -8

OpenStudy (anonymous):

|5x – 2| >= 8 Two scenarios: First one: Just drop the bars. Second one: Drop the bars, switch the inequality sign, and make the 8 negative.

OpenStudy (anonymous):

dont think it is little tricky one

OpenStudy (anonymous):

5x-2<=-8?

OpenStudy (anonymous):

haah i think your right @uzumakhi i think @melbel just knows her stuff realllly welll ((: lol

OpenStudy (anonymous):

was 5x-2<=-8 right ?@melbel

OpenStudy (anonymous):

* @melbel

OpenStudy (anonymous):

yes its right

OpenStudy (anonymous):

okay sweet what do i do next?

OpenStudy (anonymous):

add 2 to both sides and after that divide by 5

OpenStudy (anonymous):

5x<=-10 would this be right after adding 2 to both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i divide everything by 5 x<=-2 would this be right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x<=-2 would be my final anwser?

OpenStudy (anonymous):

what about second condition have you forget about it?

OpenStudy (anonymous):

yea i kinda did lol

OpenStudy (anonymous):

are you going to do second one or not?

OpenStudy (anonymous):

no i figured it out lol on paper thank you so much for help((:

OpenStudy (anonymous):

ok